y ln x ln y dx + dy = 0


1
Expert's answer
2021-09-02T07:25:47-0400
"y\ln y\ln x dx+dy=0"

"\dfrac{dy}{y\ln y}=-\ln xdx"

Integrate


"\int\dfrac{dy}{y\ln y}=-\int\ln xdx"

"\int\ln xdx=x\ln x-\int x(\dfrac{1}{x})dx=x\ln x-x-\ln C"

"\ln(|\ln y|)=-x\ln x+x+\ln C"

"\ln y=C(e^{x-x\ln x})"

"y=e^{C(e^{x-x\ln x})}"